2 cos^2 sigma sign + sin sigma symbol-2=0
interval [0,2π)

Answer by lwsshak3(11628)

2 cos^2 sigma sign + sin sigma symbol-2=0
interval [0,2π)
***
use x for sigma:
solve for x
2cos^x+six-2=0
2(1-sin^2x)+sinx-2=0
2-2sin^2x+sinx-2=0
-2sin^2x+sinx=0
sinx(-2sinx+1)=0
..
sinx=0
x=0
or
-2sinx+1=0
sinx=1/2
x=π/4, 3π/4